MNO, Inc., a publicly traded manufacturing firm in the United States, has provided the following financial information in its application for a loan. The market value of equity is 1.2059 times the book value of debt. Retained earnings are 3.17% of total assets. Sales are 106.67% of total assets. Earnings before interest and taxes are 46.67% of total assets. Finally, Working capital is 25.33% of total assets.

a) What is the Altman discriminant function value for MNO, Inc.? Recall that:

b) Should you approve MNO, Inc.'s application to your bank for a $500,000 capital expansion loan?

c) If sales for MNO dropped by 68.75%, the market value of equity was only half of the book value of debt, and the earnings before interest and taxes were only 10% of total assets, but everything else was unchanged would your credit decision change?
Answer #1

Altman’s discriminant function is given by: Z = 1.2X1 + 1.4X2 + 3.3X3 + 0.6X4 + 1.0X5

where X1 = Working capital/total assets (TA) = 25.33% = 0.2533

X2 = Retained earnings/TA = 3.17% = 0.0317

X3 = EBIT/TA = 46.67% = 0.4667

X4 = Market value of equity/long term debt = 1.2059

X5 = Sales/TA = 106.67% = 1.0667

Part (a)

the Altman discriminant function value for MNO, Inc. = 1.2 x 0.2533 + 1.4 x 0.0317 + 3.3 x 0.4667 + 0.6 x 1.2059 + 1 x 1.0667 = 3.67869

Part (b)

Since the Z-score of 3.67869 which is greater than 2.99, MNO Inc.’s application for a capital expansion loan should be approved.

Part (c)

If sales for MNO dropped by 68.75%, the market value of equity was only half of the book value of debt, and the earnings before interest and taxes were only 10% of total assets,

X3 = EBIT/TA = 10% = 0.10

X4 = Market value of equity/long term debt = 0.5

X5 = Sales/TA = 106.67% - 68.75% = 0.3792

the Altman discriminant function value for MNO, Inc. = 1.2 x 0.2533 + 1.4 x 0.0317 + 3.3 x 0.10 + 0.6 x 0.5 + 1 x 0.3792 = 1.35754

Since the Z-score of 1.35754 which is lower than 2.99, there will be change in decision. MNO Inc.’s application for a capital expansion loan should not be approved.